The perimeter of a rectangle is 34 feet. Find the width of the rectangle if the length is 5 feet more than twice the width.

Answer by Deina(147)


2L + 2W = 34
l = 2w + 5

2(2w + 5) + 2w = 34
4w + 10 + 2w = 34 Combine like terms:
6w + 10 = 34

Subtract 10 from each side:
6w = 24

Divide both sides by 6
w = 4 Width equals 4

Substitute 4 for w calculate length:
l = 2(4) +5
l = 13 Length = 13

replace l with 13 and w with 4 to check your answer:
2(13) + 2(4) = 34 and bob's your uncle!